OpenStudy (anonymous):

Given the two expressions shown below: A.64 square root plus 5 square root B.64 square root plus 4 square root Which statement best describes the two expressions?

OpenStudy (anonymous):

Both are rational. Both are irrational. A is rational, but B is irrational. A is irrational, but B is rational.

OpenStudy (anonymous):

\[\sqrt{64}=?,\sqrt{4}=?,\sqrt{5}=?\] if the result is an integer ,then rational,otherwise not.

OpenStudy (anonymous):

is that 64 symbol plus the 5 symbol

OpenStudy (anonymous):

so its b

OpenStudy (anonymous):

A. \[\sqrt{64}+\sqrt{5}=8+\sqrt{5}\] B.

OpenStudy (anonymous):

B. \[\sqrt{64}+\sqrt{4}=8+2=10\]

OpenStudy (anonymous):

so is it b

OpenStudy (anonymous):

no, A is irrational as square root 5 cannot be represented as an integer. B is rational i have shown above.